OpenStudy (anonymous):

Compare and contrast the balance sheet treatment of a car purchased with cash and a car purchased on credit. How is equilibrium of the basic accounting equation maintained in both instances? Clarify why the methods may be treated differently

OpenStudy (anonymous):

Gosh, Im so stupid with this stuff. Can someone point me in the right direction and help me get started on this?

OpenStudy (amistre64):

The equilibrium of the accounting equation simply means that: Total Assets = Total Liabilities + Equity (the stuff you can claim as your own after your debts are paid off.) The equation maintains and records all transactions that bring money in, take money out, or move the money around. Lets start out with some account balances to see what happens to the balance sheet as we make transations: [Balance Sheet] Asset ------- Cash 3000 Car 0 ==== Total Assets .......................$ 3000 Liabilities -------- Accnts Payable 2000 Notes Payable 500 ==== Total Liabilities.......................$ 2500 Equity ------ Capitol Stock 500 ==== Total Equity ...........................$ 500 3000 = 2500 + 500 ; Assets = Liabilities + Equity ............................................................................... The balance sheet treatment of a car purchased with cash results in the amount of Cash in the Cash account is decreased by the amount you paid for the car. The first entry here in my mock ledger is an X, which identifies our account as an (A)sset, (L)iability, or (E)quity X | Account | Debit | Credit --------------------------- A | Car | 1500 | | --------------------------- A | Cash | | 1500 | --------------------------- The Accounting equation is affected only by money moving around between one asset to another, and the equilibrium of the equation remains intact since nothing is gained or lost. The (Asset) Car Account increases by 1500, and the (Asset) Cash Account decreases by 1500. By recording it in more than one account allows us to see HOW the money moved. [Balance Sheet] Asset ------- Cash 1500 Car 1500 ==== Total Assets .......................$ 3000 Liabilities -------- Accnts Payable 2000 Notes Payable 500 ==== Total Liabilities.......................$ 2500 Equity ------ Capitol Stock 500 ==== Total Equity ...........................$ 500 3000 = 2500 + 500 ; Assets = Liabilities + Equity ............................................................................ Now lets see it bought on credit. X | Account | Debit | Credit -------------------------------- A | Car | 1500 | | -------------------------------- L | Acct Payable | | 1500 | -------------------------------- Different Accounts are affected, but the equilibrium of the accounting equation is maintained becasue we kept track of which accounts are affected by the transaction. [Balance Sheet] Asset ------- Cash 3000 Car 1500 ==== Total Assets .......................$ 4500 Liabilities -------- Accnts Payable 3500 Notes Payable 500 ==== Total Liabilities.......................$ 4000 Equity ------ Capitol Stock 500 ==== Total Equity ...........................$ 500 4500 = 4000 + 500 ; Assets = Liabilities + Equity ........................................................................... Thats as good as I can answer :) Hope it helps..

OpenStudy (anonymous):

do you work in this field or something?

OpenStudy (amistre64):

Nah, just recalling what I took from accounting in college ..

OpenStudy (anonymous):

hi friend. I really need some more help and you seem to be the only one I understand. I have a team assignment due on Friday. But I only have 1 little section of it. The companys we are using are Barnes and Noble and Netflix. My part of the assignment is the last bullet point. I dont need you to do it for me, but I need you to explain it for me as I have no idea how to do this. I dont even know what it means...Below is a break down of the assignment, and again, its just the last bullet point I need to focus on. I need a lifesaver right about now...any assistance? Resource: Review of Financial Statements Paper Prepare a summary comparing the organizations’ two most recent fiscal years based on the following: • Discuss the interrelationships noted among data provided by each statement. • Describe where key components of the basic accounting equation are illustrated in the organization’s financial statements. • Describe any control techniques or issues discussed in the statements.